Couldn't figure out the answer to this on my 5th graders math assignment:
6+(3/5x)=x, can you help?

Answer by queenofit(120)

First always put your X on one side.
I find it easier to think of the numbers and the variables as opposing teams. You want all the numbers on their teams side and all the variables on the other.
so,
when you move something to the other side of the equal sign, you have to change its sign.. If it is positive on one side, when you move it across the equal it becomes negative and vice versa.
so,
6= -3/5x + x remember there is a 1 in front of your x. you will have
6= -3/5x + 5/5x
6= 2/5x
you have to get rid of the 5 by multiplying 5 on each side of the = sign
6 * 5 = 2/5x * 5 the fives on the right cancell out because in 2/5, 2 is being divided by 5 so you are left with 2x on the right of the =.
6 * 5 is 30 so you have the new problem of
30=2x to get rid of the 2 in front of the x you must divide each side by 2.
15=x
Remember to if you do something on one side of the equal sign you must do the same on the other side of the equal sign...